# Chain-of-Factors Paper-Reviewer Matching

### v2

@@ 4.4. Effectiveness in Each Factor
One may suspect that some baselines are more powerful than CoF in a certain factor, but finally underperform CoF in paper-reviewer matching just because CoF takes more factors into account and adopts chain-of-factors matching. To dispel such misgivings, we examine the performance of compared methods in three tasks – semantic retrieval, topic classification, and citation prediction – corresponding to the three factors, respectively. Specifically, for each submission paper $p$ in the KDD dataset [^3] We use the KDD dataset for experiments in Sections 4.4 and 4.5 because the required information of each paper (e.g., venue, year, references, fields) is stored when we construct the dataset. By contrast, such information is largely missing in the other three datasets., we sample 100 candidates among which only one is “relevant” to $p$ and the other 99 are “irrelevant”. Here, the meaning of “relevant” depends on the examined task. For topic classification, we sample one of $p$’s fields, and the “relevant” candidate is the name of that field; the “irrelevant” candidates are names of other randomly sampled fields. For citation prediction, we select one of the papers cited by $p$ as the “relevant” candidate; the “irrelevant” candidates are chosen from candidate reviewers’ previous papers not cited by $p$. For semantic retrieval, we follow (Jin et al., 2019) to conduct a title-to-abstract retrieval task, where the query is $p$’s title, the “relevant” candidate is $p$’s abstract, and the “irrelevant” candidates are sampled from other papers’ abstracts. Note that for the topic classification task, the instructions used by CoF for paper-reviewer matching are no longer suitable, so we adopt a new instruction “Tag a scientific paper with relevant scientific topic classes.”.
@@ 5. Related Work
Paper-Reviewer Matching. Following the logic of the entire paper, we divide previous paper-reviewer matching methods according to the factor they consider. In earlier times, semantic-based approaches use bag-of-words representations, such as tf–idf vectors (Yarowsky and Florian, 1999; Hettich and Pazzani, 2006; Charlin and Zemel, 2013) and keywords (Tang and Zhang, 2008; Protasiewicz, 2014) to describe each submission paper and each reviewer. As a key technique in classical information retrieval, probabilistic language models have also been utilized in expert finding (Balog et al., 2006). More recent semantic-based methods have started to employ context-free word embeddings (Zhao et al., 2018; Zhang et al., 2020) for representation. Topic-based approaches leverage topic models such as Latent Semantic Indexing (Dumais and Nielsen, 1992; Li and Hou, 2016), Probabilistic Latent Semantics Analysis (Karimzadehgan et al., 2008; Conry et al., 2009), and Latent Dirichlet Allocation (and its variants) (Liu et al., 2014; Kou et al., 2015; Jin et al., 2017) to infer each paper/reviewer’s topic distribution. This idea is recently improved by exploiting embedding-enhanced topic models (Qian et al., 2018; Anjum et al., 2019). Inspired by the superiority of contextualized language models to context-free representations, recent studies (Singh et al., 2023; Stelmakh et al., 2023) apply scientific PLMs such as SPECTER (Cohan et al., 2020), SciNCL (Ostendorff et al., 2022), and SPECTER 2.0 (Singh et al., 2023) to perform paper-reviewer matching. These PLMs are pre-trained on a large amount of citation information between papers. For a more complete discussion of paper-reviewer matching studies, one can refer to a recent survey (Zhao et al., 2022). Note that most of the aforementioned approaches take only one factor into account, resulting in an incomprehensive estimation of the paper-reviewer relevance. In comparison, CoF jointly considers the semantic, topic, and citation factors with a unified model.
@@
6. Conclusions
In this work, we present a Chain-of-Factors framework that jointly considers semantic, topic, and citation signals in a step-by-step, coarse-to-fine manner for paper-reviewer matching. We propose an instruction-guided paper encoding process to learn factor-aware text representations so as to model paper-reviewer relevance of different factors. Such a process is facilitated by pre-training an instruction encoder and a paper encoder with a contextualized language model backbone. Experimental results validate the efficacy of our CoF framework on four datasets across various fields. Further ablation studies explain the major reasons why CoF is superior to the baselines: (1) CoF comprehensively considers three factors rather than just one, (2) CoF chains three factors to facilitate step-by-step, coarse-to-fine relevant paper selection instead of merging all factors in one step, and (3) CoF improves upon the baselines in each of the factors. We also conduct analyses on how the composition of each reviewer’s publication profile will affect the paper-reviewer matching performance.
@@ A.1.2. Construction of the KDD Dataset
We rely on the Microsoft Academic Graph (MAG) (Sinha et al., 2015) to extract each paper’s title, abstract, venue, and author(s). The latest KDD conference available in our downloaded MAG is KDD 2020. Therefore, we first retrieve all KDD 2020 papers from MAG as potential “submission” papers. Then, we select those researchers meeting the following two criteria as candidate reviewers: (1) having published at least 1 KDD paper during 2018-2020, and (2) having published at least 3 papers in “top conferences”. Consistent with the definition in Section 4.5, “top-conferences” refer to the 75 conferences listed on CSRankings in 2020, including KDD. Guided by our observations in Section 4.5, for each candidate reviewer $r$, we include all of its papers published in 2019 or earlier to form its publication profile $\mathcal{Q}_{r}$. Next, we randomly sample about 200 papers from KDD 2020. For each sampled paper, we select 20 candidate reviewers for annotation. We do our best to ensure that conflict-of-interest reviewers (e.g., authors and their previous collaborators) are not selected. To reduce the possibility that none of the selected reviewers is relevant to the paper (which makes the paper useless in evaluation), reviewers sharing a higher TPMS score (Charlin and Zemel, 2013) with the paper are more likely to be selected for annotation. Finally, we invite 5 annotators to independently rate each pair of (paper, selected reviewer) according to the “0”-“3” relevance scheme. The final score between a paper and a reviewer is the average rating from the annotators rounded to the nearest integer. We remove papers that: (1) do not have any selected reviewer with an annotated relevance score greater than or equal to “2” or (2) annotators are not able to judge its relevance to some candidate reviewers, resulting in 174 papers in the final dataset. On average, each paper in our KDD dataset has 2.10 reviewers with a relevance score of “3”, 3.05 reviewers with a score of “2”, 6.32 reviewers with a score of “1”, and 8.53 reviewers with a score of “0”.

### v3

@@ 4.4. Effectiveness in Each Factor
One may suspect that some baselines are more powerful than CoF in a certain factor, but finally underperform CoF in paper-reviewer matching just because CoF takes more factors into account and adopts chain-of-factors matching. To dispel such misgivings, we examine the performance of compared methods in three tasks – semantic retrieval, topic classification, and citation prediction – corresponding to the three factors, respectively. Specifically, for each submission paper $p$ in the KDD dataset [^2] We use the KDD dataset for experiments in Sections 4.4 and 4.5 because the required information of each paper (e.g., venue, year, references, fields) is stored when we construct the dataset. By contrast, such information is largely missing in the other three datasets., we sample 100 candidates among which only one is “relevant” to $p$ and the other 99 are “irrelevant”. Here, the meaning of “relevant” depends on the examined task. For topic classification, we sample one of $p$’s fields, and the “relevant” candidate is the name of that field; the “irrelevant” candidates are names of other randomly sampled fields. For citation prediction, we select one of the papers cited by $p$ as the “relevant” candidate; the “irrelevant” candidates are chosen from candidate reviewers’ previous papers not cited by $p$. For semantic retrieval, we conduct a title-to-abstract retrieval task, where the query is $p$’s title, the “relevant” candidate is $p$’s abstract, and the “irrelevant” candidates are sampled from other papers’ abstracts. Note that for the topic classification task, the instructions used by CoF for paper-reviewer matching are no longer suitable, so we adopt a new instruction “Tag a scientific paper with relevant scientific topic classes.”.
@@ 5. Related Work
Paper-Reviewer Matching. Following the logic of the entire paper, we divide previous paper-reviewer matching methods according to the factor they consider. In earlier times, semantic-based approaches use bag-of-words representations, such as tf–idf vectors (Yarowsky and Florian, 1999; Hettich and Pazzani, 2006; Charlin and Zemel, 2013) and keywords (Tang and Zhang, 2008; Protasiewicz, 2014) to describe each submission paper and each reviewer. As a key technique in classical information retrieval, probabilistic language models have also been utilized in expert finding (Balog et al., 2006). More recent semantic-based methods have started to employ context-free word embeddings (Zhao et al., 2018; Zhang et al., 2020) for representation. Topic-based approaches leverage topic models such as Latent Semantic Indexing (Dumais and Nielsen, 1992; Li and Hou, 2016), Probabilistic Latent Semantics Analysis (Karimzadehgan et al., 2008; Conry et al., 2009), and Latent Dirichlet Allocation (and its variants) (Liu et al., 2014; Kou et al., 2015; Jin et al., 2017) to infer each paper/reviewer’s topic distribution. This idea is recently improved by exploiting embedding-enhanced topic models (Qian et al., 2018; Anjum et al., 2019). Inspired by the superiority of contextualized language models to context-free representations, recent studies (Singh et al., 2023; Stelmakh et al., 2023) apply scientific PLMs (Zhang et al., 2024) such as SPECTER (Cohan et al., 2020), SciNCL (Ostendorff et al., 2022), and SPECTER 2.0 (Singh et al., 2023) to perform paper-reviewer matching. These PLMs are pre-trained on a large amount of citation information between papers. For a more complete discussion of paper-reviewer matching studies, one can refer to a recent survey (Zhao et al., 2022). Note that most of the aforementioned approaches take only one factor into account, resulting in an incomprehensive estimation of the paper-reviewer relevance. In comparison, CoF jointly considers the semantic, topic, and citation factors with a unified model.
@@
6. Conclusions and Future Work
In this work, we present a Chain-of-Factors framework that jointly considers semantic, topic, and citation signals in a step-by-step, coarse-to-fine manner for paper-reviewer matching. We propose an instruction-guided paper encoding process to learn factor-aware text representations so as to model paper-reviewer relevance of different factors. Such a process is facilitated by pre-training an instruction encoder and a paper encoder with a contextualized language model backbone. Experimental results validate the efficacy of our CoF framework on four datasets across various fields. Ablation studies reveal the key reasons behind the superiority of CoF over the baselines: (1) CoF takes into account three factors holistically rather than focusing on just one, (2) CoF integrates these three factors in a progressive manner for relevant paper selection, rather than combining them in a single step, and (3) CoF outperforms the baselines across each individual factor. We also conduct analyses on how the composition of each reviewer’s publication profile will affect the paper-reviewer matching performance.
As for future work, we strongly believe that deploying our model to real conference management systems would largely increase the practical value of this paper. Also, it would be interesting to generalize our chain-of-factors matching framework to other tasks that aim to learn the proximity between two text units.
Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
@@ A.1.1. Pre-training Data
Hard Negatives. Cohan et al. (Cohan et al., 2020) show that a combination of easy negatives and hard negatives boosts the performance of their contrastive learning model. Following their idea, given a factor $\phi$ and its training sample $(p,q^{+},q^{-}_{1},q^{-}_{2},...,q^{-}_{T})$, we take in-batch negatives (Karpukhin et al., 2020) as easy negatives and adopt the following strategies to find hard negatives: when $\phi={\rm semantic}$, $q^{-}_{t}$ is a hard negative if it is shown to users but not clicked given the query $p$; when $\phi={\rm topic}$, $q^{-}_{t}$ is a hard negative if it shares the same venue but does not share any fine-grained field with $p$; when $\phi={\rm citation}$, $q^{-}_{t}$ is a hard negative if $q^{+}$ cites $q^{-}_{t}$ but $p$ does not cite $q^{-}_{t}$.
@@ A.1.2. Construction of the KDD Dataset
We rely on the Microsoft Academic Graph (MAG) (Sinha et al., 2015) to extract each paper’s title, abstract, venue, and author(s). The latest KDD conference available in our downloaded MAG is KDD 2020. Therefore, we first retrieve all KDD 2020 papers from MAG as potential “submission” papers. Then, we select those researchers meeting the following two criteria as candidate reviewers: (1) having published at least 1 KDD paper during 2018-2020, and (2) having published at least 3 papers in “top conferences”. Consistent with the definition in Section 4.5, “top-conferences” refer to the 75 conferences listed on CSRankings in 2020, including KDD. Guided by our observations in Section 4.5, for each candidate reviewer $r$, we include all of its papers published in 2019 or earlier to form its publication profile $\mathcal{Q}_{r}$. Next, we randomly sample about 200 papers from KDD 2020. For each sampled paper, we select 20 candidate reviewers for annotation. We do our best to ensure that conflict-of-interest reviewers (e.g., authors and their previous collaborators) are not selected. To reduce the possibility that none of the selected reviewers is relevant to the paper (which makes the paper useless in evaluation), reviewers sharing a higher TPMS score (Charlin and Zemel, 2013) with the paper are more likely to be selected for annotation. Finally, we invite 5 annotators to independently rate each pair of (paper, selected reviewer) according to the “0”-“3” relevance scheme. During this process, we provide each annotator with the paper title, the paper abstract, the reviewer’s name, and the reviewer’s previous papers (sorted by their citation counts from high to low). The final score between a paper and a reviewer is the average rating from the annotators rounded to the nearest integer. We remove papers that: (1) do not have any selected reviewer with an annotated relevance score greater than or equal to “2” or (2) annotators are not able to judge its relevance to some candidate reviewers, resulting in 174 papers in the final dataset. On average, each paper in our KDD dataset has 2.10 reviewers with a relevance score of “3”, 3.05 reviewers with a score of “2”, 6.32 reviewers with a score of “1”, and 8.53 reviewers with a score of “0”.